# Noise-resilient host for electron qubit operation up to 0.4 K

### v1

@@
Noise-resilient host for electron qubit operation up to 0.4 K
@@
(February 3, 2025)
Cryogenic solid neon has recently emerged as a pristine solid-state material platform for hosting single electron qubits. At temperatures of $\sim$10  mK, electron-on-solid-neon (eNe) charge qubits have shown exceptionally long coherence times and high single-qubit gate fidelities. Further expanding the capabilities of this platform to multiple qubits in a scalable fashion will require careful understanding and the corresponding optimization of the underlying noise environment. Here, we demonstrate the remarkable robustness enabled by the neon layer in eNe charge qubits against charge and thermal noise by operating them away from the charge-insensitive sweet-spot and at elevated temperatures. Without optimization, the measured charge (voltage) noise spectral density on neon is comparable to the best in conventional gate-defined semiconductor quantum dots. Up to 0.4 K, the qubit coherence time with dynamical decoupling can still reach the relaxation limit of $\sim$1 microsecond. These observations highlight the potential of solid neon to facilitate a quiet charge environment and eNe qubits’ potential for higher-temperature operation in a scaled-up system.
@@
Solid-state electron qubits are inherently affected by decoherence mechanisms in their host materials. Spectral characterization has revealed a typical $1/f$ noise distribution for both charge and spin qubits, often attributed to individual charge fluctuators at the surfaces and interfaces of semiconductor or superconducting materials [1, 2, 3]. Extensive effort has been made to extend electron qubit coherence times by reducing the noise density in the surrounding environment and minimizing the sensitivity of qubits to noise  [4, 2, 5]. Recently, we demonstrated solid neon as a novel host material that traps electrons at the neon-vacuum interface [6, 7]. Using a cavity quantum electrodynamics architecture, we address the charge states of eNe by coupling them to superconducting resonators. At the charge sweet-spot, eNe qubits are first-order insensitive to charge noise, showing long coherence times $T_{2}^{*}$ of up to $\sim$50 $\mu$s (ref [7]), which is nearly four orders of magnitude longer than for reported semiconductor charge qubits [8]. In principle, this should translate quite favorably for operating eNe spin qubits, assuming that isotopically purified neon is used [9, 10, 11].
@@
There remains, however, more to be understood about the nature of eNe charge qubits (which we will simply refer to as “eNe qubits” for the remainder of the manuscript) and their environment. The specific electron trapping mechanism on solid neon is unclear, though it is likely that disorder in the neon surface plays a key role [12]. Moreover, the environmental coherence-limiting factors in current devices remain elusive. Investigating the performance of eNe qubits away from the sweet-spot and at elevated temperatures, when the qubit is subject to charge and thermal noise, can provide valuable insight on the coupling between eNe and the environment, revealing the role of the neon host and paving the way for improved qubit performance and stability. This is crucial for scaling up the eNe qubit platform, as environmental noise acts as a major obstacle to the precise creation and consistent retention of entangled multi-qubit states [1, 13]. It is also important for understanding the limitations of spin-state control [2, 14, 9] of eNe using electrically sensitive mechanisms such as synthetic spin-orbit coupling [15, 16, 17, 18, 19] or exchange interactions [20, 21, 22, 23]. Furthermore, operating qubits at elevated temperatures can mitigate many engineering constraints due to limited cooling power at base temperature, advantageous for scaling [24, 25, 26].
@@
Fig. 1. : eNe charge qubit coupled to a TiN high-impedance superconducting resonator. a, Illustration of the high impedance TiN superconducting resonator with two identical electron traps. Microwave (MW) signals to control and read out the electron qubit’s charge state are sent through the resonator via the input and output ports. Gate electrodes are patterned with on-chip low-pass filters to apply DC bias voltage to the qubits while maintaining the resonator’s quality factor. The metal plane between the resonator pins is connected to the ground plane via aluminum wire bonds. b, False-color scanning electron micrograph image of the electron trapping area on the right end of the resonator with trap gates (RTG). c, Cross-section schematic of the electron trapping area.
@@
Fig. 2. : Spectroscopic and coherence properties of an eNe charge qubit. a, Normalized microwave transmission amplitude ($A/A_{0}$) centered around the resonator frequency versus the relative resonator bias voltage $\Delta V_{\text{res}}$ as described in (b). Two avoided crossings appear when the eNe qubit comes into resonance with the resonator. b, Two-tone qubit spectroscopy measurement displaying the transmission phase response at the resonator frequency $\omega_{\text{r}}$ versus $\Delta V_{\text{res}}$, for different weak probe tones concurrently sent in at $f_{\text{drive}}$. c, Ramsey fringes at the charge sweet-spot, marked with the red arrow in (b), with fitted $T_{2}^{*}$ of 8.2 µs. $P_{e}$ is the qubit’s excited-state population. d, $T_{1}$ = 11.6 µs and $T_{2}^{\text{echo}}$ = 21.6 µs at the charge sweet-spot.
@@
Our device consists of a split superconducting resonator made of a 30 nm thick TiN film grown on a $\langle 111\rangle$-oriented intrinsic silicon (Si) substrate by atomic layer deposition [27], as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka. An electron trap is positioned at each end of the resonator, with the broader goal of coupling two distant eNe qubits via the resonator bus [28]. Each electron trap features a simplified rectangular structure compared to the previously used oval-shaped design [6, 7]. Around the traps, the Si substrate is etched down by approximately 250 nm to host the thin neon layer, as illustrated in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb. In this work, we utilize the differential mode of the resonator to couple to motional states of electrons trapped on solid neon (see Supplementary I), with the microwave electrical field pointing from one resonator pin to the other [29], as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kc. Considering the first two charge states of eNe, the coupled system can be described by the Jaynes-Cummings Hamiltonian [14]:
@@
As an improvement from previous work, we leverage the high kinetic inductance ($\sim$20 pH/$\Box$) of the thin TiN film to enhance the qubit-resonator coupling strength [30]. The estimated equivalent lumped element impedance for the differential mode is $Z_{r}\sim 600\,\Omega$ (see Supplementary I), approximately ten times that of the previous niobium (Nb) device [6, 29]. Since $g\propto\omega_{r}\sqrt{Z_{r}}$, we expect the high impedance resonator to support a coupling strength achieving $g/2\pi\sim 10\,\text{MHz}$ level [30]. Meanwhile, several DC gates are placed around the electron trapping area to tune the qubit transition frequency, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka-b. In order to minimize microwave leakage of the resonator mode through the DC gates, all gates are equipped with on-chip low-pass filters with a $\sim$0.5 GHz cutoff frequency providing over 60 dB attenuation at resonator frequency. This is particularly important in our device given that for high-impedance resonators, the parasitic gate capacitance can be comparable to the resonator capacitance [31, 32]. With this design, the resonator maintains a narrow linewidth $\kappa/2\pi$ of 0.38 MHz (see Supplementary I). Before performing electron experiments, a thin layer of neon was grown on the surface of the device following the same procedure as in our previous work [7], resulting in less than 1 MHz redshift of the resonator frequency. Electrons are emitted from a tungsten filament and bound to the surface of the neon film. Only when both the neon film is present on the sample surface and electrons are emitted do we see signatures of trapped eNe qubits strongly coupled to the resonator (see Supplementary II).
Fig. 3. : Dephasing of eNe charge qubit Q1. a, Calculated $1/T_{e}^{\rm{CPMG}}$ for various qubit bias points and refocusing pulse numbers $N$. With the increase of $N$, the $1/T_{e}^{\rm{CPMG}}$ of the qubit biased near the charge sweet-spot approaches the limit of $2T_{1}$. b, The pure dephasing time $T_{\phi}$ increases as a function of $N$ when the qubit biased away from charge sweet-spot, with a power-law fit of $T_{\phi}\propto N^{0.61}$. c, Repeated Ramsey fringes measured near qubit sweet-spot for 128 iterations, with each record taking 33 s. d, Detuning $\Delta_{dq}$ between drive tone and qubit frequency during the Ramsey measurements, revealing stochastic frequency shifts.
@@
As illustrated in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kc, the local neon profile and the trap structure define the potential energy landscape seen by the electrons. Recent theoretical work emphasized the important role that the neon surface plays in defining the qubit’s Hamiltonian [12], whose exact form remains unclear. Here we approximate each eNe qubit’s transition frequency with a generalized hyperbolic model, capturing the measured qubit spectroscopic features:
@@
Bringing eNe qubit’s transition frequency $\omega_{\rm{q}}$ onto resonance with $\omega_{\rm{r}}$ results in the vacuum Rabi splitting in the resonator’s transmission spectrum. An example of this is shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka for Q1, one of the three qubits we characterized, with the electron-photon coupling strength $g/2\pi=6.43$ MHz and on-resonance qubit linewidth $\gamma/2\pi=3.81$ MHz (see Supplementary III-VI for details of characterizations on the three qubits Q1-3). Compared to the previous Nb resonator [6, 7], the higher impedance of the TiN resonator enhances the qubit-resonator coupling strength, with a maximum observed $g/2\pi$ of approximately $16$ MHz (see Q3 in Supplementary VI). We further mapped the qubit spectrum by applying a second drive tone and probing at the resonator frequency at low power, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb. The extracted qubit frequency follows a hyperbolic dependence on the DC voltage $V_{\text{res}}$ applied to the resonator with a charge sweet-spot at 5.065 GHz (see Supplementary IV). The sweet-spot frequency, coupling strength, and sensitivity to DC biases ($\partial f_{\text{q}}/\partial V_{\text{res}}$) vary between qubits (see Supplementary III). This variation suggests some randomness in the local trapping potential, which determines the qubit’s minimum transition frequency and the projection of the electron’s dipole moment along the resonator mode and the applied DC field.
@@
Next, we characterize the coherence performance of the qubit biased at its charge sweet-spot. Figure Noise-resilient host for electron qubit operation up to 0.4 Kc-d show the measured $T_{1}$, $T_{2}^{*}$, and $T_{2}^{\text{echo}}$ of 11.6 µs, 8.2 µs, and 22.6 µs, respectively. Unlike the qubit reported in our previous works [7], the relaxation time ($T_{1}$) of eNe qubits observed on the new TiN device is generally not Purcell-limited at their charge sweet-spots, with over 100 MHz detuning from the resonator ($\Delta_{\rm{rq}}/2\pi$). The Purcell rate of Q1 to resonator mode is $\Gamma_{\rm{r}}\,=\,\kappa g^{2}/\Delta_{\rm{rq}}^{2}\,=\,1/3.9\,\text{ms}^{-1}$, which indicates that non-radiative decay channels dominate the energy relaxation of Q1. Measurements of the Q1’s $T_{1}$ at various bias frequencies (see Noise Spectroscopy section) further confirm that the non-radiative decay dominates the energy relaxation unless the qubit frequency is tuned much closer to the resonator frequency. Additionally, the fact that the $T_{2}^{\rm{echo}}$ approaches $2T_{1}$ indicates that the quasi-static noise is the dominating dephasing factor for Q1 at its charge sweet-spot [37, 38].
@@
Fig. 4. : Noise spectroscopy of qubit Q1. Main plot between 0.01 $\sim$ 1 MHz: Colored dots: Total noise density derived from dynamical decoupling data at different qubit bias points. Main plot between $10^{-3}$ to $10^{-1}$ Hz: Blue dots: Calculated total noise density from long-term Ramsey measurements when biased at charge sweet-spot. Main plot near 5 GHz: Green triangles: Transverse noise of the eNe qubit. Inset plot i: Equivalent charge (voltage) noise on the resonator electrode calculated. Inset plot ii: Zoom-in of transverse noise. Gray dashed lines: power-law fits of frequency dependent noise.
@@
Fluctuations in the charge environment can cause stochastic frequency shifts of the eNe qubit, which leads to qubit dephasing. Armed with high fidelity control and a known voltage (charge) lever arm, we first characterize the high-frequency noise spectral density of Q1 using the Carr-Purcell-Meiboom-Gill (CPMG) sequence with $N=0,1,2,4,6,8,12$, and 16 refocusing pulses. All bias points had positive $\Delta V_{\rm{res}}$ as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb, with a maximum sensitivity of $\partial f_{\rm{q}}/\partial V_{\rm{res}}=180.7$ MHz/mV, corresponding to a qubit frequency detune of 15.9 MHz from the charge sweet-spot. The measured average energy relaxation time at all bias points is 11.94 $\pm$ 0.3 µs, indicating that the same non-radiative decay channel dominates. We fit the measured coherence decay to extract the decoherence time $T_{2}^{\text{CPMG}}$ and pure dephasing time $T_{\phi}$ (See Methods). As shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka, in the absence of refocusing pulses (i.e., the Ramsey measurement), the decoherence rate $1/T_{2}^{\text{CPMG}}$ increases with sensitivity to charge noise. At $\partial f_{\rm{q}}/\partial V_{\rm{res}}=180.7$ MHz/mV, $T_{2}^{*}$ decreases to 1.93 µs. Introducing refocusing pulses enhances $T^{\text{CPMG}}_{2}$ at every biasing point, approaching two times of $T_{1}$ as $N$ increased. At the charge sweet-spot, the rapid saturation of $T_{2}^{\text{CPMG}}$ with increasing $N$ indicates a low noise power density in the relatively high frequency range (100 kHz). When the qubit was biased at a point more sensitive to charge noise, more pulses were required to extend the $T_{2}^{\text{CPMG}}$ towards the energy relaxation limit. In Supplementary V and VI, we show that for the other two eNe qubits, a single refocusing pulse is insufficient to mitigate the majority of noises at the qubits’ sweet-spot, suggesting higher noise density at the high-frequency range. This variation in noise behavior suggests some complexities of the local noise environment experienced by individual electrons on the neon surface.
@@
To benchmark different electron qubit material platforms, we compare the environmental charge noise measured as voltage fluctuations on adjacent gate electrodes, situated at a distance on the scale of around one hundred nanometers from the qubits. Specifically, we examine electrons trapped on solid neon and those confined in semiconductor materials, which are the leading platforms for hosting electron qubits. Table 1 summarizes our results on eNe and compares them with data from literature on semiconductor qubits. Because of the clean qubit environment provided by the neon interface, the charge noise experienced by eNe is orders of magnitude lower than that in typical semiconductor materials, approaching some of the best performance. Since the susceptibility of electron spin qubit to charge noise through effects including spin-orbit coupling and exchange interaction is in general less than that of electron charge qubit, as shown in the “$\partial f_{q}/\partial V$” column of Tab. Noise-resilient host for electron qubit operation up to 0.4 K, we foresee that electron spin qubit on solid neon could have even better coherence. Further development of the electron trap and the suppression of the influence from redundant electrons on neon will further improve the performance of eNe qubits.
@@
After investigating high-frequency noise, we turn to the low-frequency noise at the charge sweet-spot. Figure Noise-resilient host for electron qubit operation up to 0.4 Kc-d show variations in Q1’s frequency revealed by performing repeated Ramsey measurements for approximately one hour [21, 20] since initially biased near its charge sweet-spot. The qubit undergoes a discrete frequency transition of varying magnitudes, as seen near the 25 minutes of the one-hour tracing, revealed by a Fast Fourier transform (FFT) of the Ramsey signals (See Methods and Supplementary IV). This suggests slow mechanisms driving qubit decoherence, akin to low-frequency dynamics in semiconductor and superconducting qubits [21, 20, 40], which can be mitigated with feedback control [41, 42]. Using a periodogram method, we convert this one-hour qubit frequency measurement into a frequency noise spectrum over $10^{-3}$ to $10^{-1}$ Hz, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 K. A power-law fit of the data gives the relation of $S\propto f^{-1.11}$. Despite the qubit is first-order insensitive to small charge (voltage) fluctuation at the sweet-spot, it can still experience slow frequency drifts due to nearby charge fluctuations, as observed in both semiconductor and superconducting qubits via second-order effects [43, 44]. Other detection methods, such as single-electron charge sensing techniques [21], could complete the noise spectrum ranging from 1 Hz to $10^{4}$ Hz.
@@
To complete the discussion, we calculate the transverse noise at the qubit frequencies contributing to energy relaxation with $1/T_{1}=\pi/2\times S(2\pi f_{\rm{q}})$ (ref. [37]). The fitted total transverse noise versus frequency follows $S\propto f^{2.22}$ between 5.065 to 5.498 GHz, with a mean of $3.6\times 10^{5}$$^{2}/\text{Hz}$ Hz, as shown in the inset of Fig. Noise-resilient host for electron qubit operation up to 0.4 K.
@@
To obtain more information about the noise environment of eNe qubits, we measure the temperature dependence of Q1’s coherence ($T_{1}$, $T_{2}^{*}$, and $T_{2}^{\rm{echo}}$) at its charge sweet-spot from 10 mK to 500 mK. The energy relaxation $T_{1}$ is well described by a model that only accounts for the coupling of a two-level quantum system to a bosonic thermal bath within the experiment’s temperature range, $T_{1}\propto\text{tanh}(\hbar\omega_{\rm{q}}/2k_{\text{B}}T)$ (Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka) (ref. [45, 46]). Similarly, the thermal population follows that of a Maxwell-Boltzmann distribution [47] (See Supplementary IV). The corresponding single electron temperature closely tracks the mixing-chamber (MXC) temperature between 100 to 500 mK and saturates around 40 mK, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka. This effective cooling of electron qubits below 100 mK is difficult to achieve with electrons in semiconductors, due to the suppressed electron-environment interaction at low temperatures [48].

### v2

# Noise-resilient solid host for electron qubits above 100 mK

@@
Noise-resilient solid host for electron qubits above 100 mK
@@
(February 18, 2025)
Cryogenic solid neon has recently emerged as a pristine solid host for single electron qubits. At $\sim$10 mK temperatures, electron-on-solid-neon (eNe) charge qubits have exhibited exceptionally long coherence times and high operation fidelities. To advance this platform towards a scalable quantum information architecture, systematic characterization of its noise feature is imperative. Here, we show the remarkable resilience of solid neon against charge and thermal noises when eNe qubits are operated away from the charge-insensitive sweet-spot and at elevated temperatures. Without optimizing neon growth, the measured charge (voltage) noise on solid neon is already orders of magnitude lower than that in most stringently grown semiconductors, rivaling the best records to date. Up to 400 mK, the eNe charge qubits operated at $\sim$ 5 GHz can maintain their echo coherence times over 1 microsecond. These observations highlight solid neon as an ideal host for quantum information processing at higher temperatures and larger scales.
@@
Solid-state electron qubits are inherently affected by decoherence mechanisms in their host materials. Spectral characterization has revealed a typical $1/f$ noise distribution for both charge and spin qubits, often attributed to individual charge fluctuators at the surfaces and interfaces of semiconductor or superconducting materials [1, 2, 3]. Extensive effort has been made to extend electron qubit coherence times by reducing the noise density in the surrounding environment and minimizing the sensitivity of qubits to noise  [4, 2, 5]. Recently, we demonstrated solid neon as a novel host material that traps electrons at the neon-vacuum interface [6, 7]. Using a circuit quantum electrodynamics architecture, we address the charge states of eNe by coupling them to superconducting resonators. At the charge sweet-spot, eNe qubits are first-order insensitive to charge noise, showing long coherence times $T_{2}^{*}$ of up to $\sim$50 µs (ref [7]), which is nearly four orders of magnitude longer than that of reported semiconductor charge qubits [8]. This behavior can be favorably translated to eNe spin qubits as well with even better predicted performance [9, 10, 11].
@@
There remains, however, more to be understood about the nature of eNe charge qubits (which we will refer to as “eNe qubits” for the remainder of the manuscript) and their environment. The specific electron trapping mechanism on solid neon is unclear, though it is likely that disorder in the neon surface plays a key role [12]. Moreover, the environmental coherence-limiting factors in current devices remain elusive. Investigating the performance of eNe qubits away from the sweet-spot and at elevated temperatures, when the qubit is subject to charge and thermal noise, can provide valuable insight into the coupling between eNe and the environment, revealing the role of the neon host and paving the way for improved qubit performance and stability. This is crucial for scaling up the eNe qubit platform, as environmental noise is a major obstacle to the precise creation and consistent retention of entangled multi-qubit states [1, 13]. It is also essential for understanding the limitations of spin-state control [2, 14, 9] of eNe using electrically sensitive mechanisms such as synthetic spin-orbit coupling [15, 16, 17, 18, 19] or exchange interactions [20, 21, 22, 23]. Furthermore, operating qubits at elevated temperatures can mitigate many engineering constraints due to limited cooling power at base temperature, advantageous for scaling [24, 25, 26].
@@
Fig. 1. : eNe charge qubit coupled to a TiN high-impedance superconducting resonator. a, Illustration of the high impedance TiN superconducting resonator with two identical electron traps. Microwave (MW) signals to control and read out the electron qubit’s charge state are sent through the resonator via the input and output ports. Gate electrodes are patterned with on-chip low-pass filters to apply DC bias voltage to the qubits while maintaining the resonator’s quality factor. The metal plane between the resonator pins is connected to the ground plane via aluminum wire bonds. b, False-color scanning electron micrograph image of the electron trapping area on the right side of the resonator with trap gates (RTG). c, Cross-section schematic of the electron trapping area.
@@
Fig. 2. : Spectroscopic and coherence properties of an eNe charge qubit Q1. a, Normalized microwave transmission amplitude ($A/A_{0}$) centered around the resonator frequency versus the relative resonator bias voltage $\Delta V_{\text{res}}$ as described in b. Two avoided crossings appear when the eNe qubit comes into resonance with the resonator. b, Two-tone qubit spectroscopy measurement displaying the transmission phase response at the resonator frequency $\omega_{\text{r}}$ versus $\Delta V_{\text{res}}$, for weak probe tones concurrently sent in at $f_{\text{drive}}$. c, Ramsey fringes at the charge sweet-spot, marked with the red arrow in b, with fitted $T_{2}^{*}$ of 8.2 µs. $P_{e}$ is the qubit’s excited-state population. d, Relaxation and Hahn echo measurements showing $T_{1}$ = 11.6 µs and $T_{2}^{\text{echo}}$ = 21.6 µs at the charge sweet-spot.
@@
Our device consists of a split superconducting resonator made of a 30 nm thick TiN film grown on a $\langle 111\rangle$-oriented intrinsic silicon (Si) substrate by atomic layer deposition [27], as shown in Fig. Noise-resilient solid host for electron qubits above 100 mKa. An electron trap is positioned at each end of the resonator, with the broader goal of coupling two distant eNe qubits via the resonator bus [28]. Each electron trap features a simplified rectangular structure compared to the previously used oval-shaped design [6, 7]. Around the traps, the Si substrate is etched down by approximately 250 nm to host the thin neon layer, as illustrated in Fig. Noise-resilient solid host for electron qubits above 100 mKb. In this work, we utilize the differential mode of the resonator to couple to motional states of electrons trapped on solid neon (see Supplementary I), with the microwave electrical field pointing from one resonator pin to the other [29], as shown in Fig. Noise-resilient solid host for electron qubits above 100 mKc. Considering the first two charge states of eNe, the coupled system can be described by the Jaynes-Cummings Hamiltonian [14]:
@@
As an improvement from previous work, we leverage the high kinetic inductance ($\sim$20 pH/$\Box$) of the thin TiN film to enhance the qubit-resonator coupling strength [30]. The estimated equivalent lumped element impedance for the differential mode is $Z_{r}\sim 600\,\Omega$ (see Supplementary I), approximately ten times that of the previous niobium (Nb) device [6, 29]. Since $g\propto\omega_{r}\sqrt{Z_{r}}$, we expect the high impedance resonator to support a coupling strength achieving $g/2\pi\sim 10\,\text{MHz}$ level [30]. Meanwhile, several DC gates are placed around the electron trapping area to tune the qubit transition frequency, as shown in Fig. Noise-resilient solid host for electron qubits above 100 mKa-b. In order to minimize microwave leakage of the resonator mode through the DC gates, all gates are equipped with on-chip low-pass filters with a $\sim$0.5 GHz cutoff frequency providing over 60 dB attenuation at resonator frequency. This is particularly important in our device given that for high-impedance resonators, the parasitic gate capacitance can be comparable to the resonator capacitance [31, 32]. With this design, the resonator maintains a narrow linewidth $\kappa/2\pi$ of 0.38 MHz (see Supplementary I). Before performing electron experiments, a thin layer of neon was grown on the surface of the device following the same procedure as in our previous work [7], resulting in less than 1 MHz redshift of the resonator frequency. Electrons are emitted from a tungsten filament and bound to the surface of the neon film. Only when both the neon film is present on the sample surface, and electrons are emitted do we see signatures of trapped eNe qubits strongly coupled to the resonator (see Supplementary II).
Fig. 3. : Decoherence of eNe qubit Q1. a, Calculated $1/T_{2}^{\rm{CPMG}}$ for various qubit bias points and refocusing pulse numbers $N$. With the increase of $N$, the $1/T_{2}^{\rm{CPMG}}$ of the qubit biased near the charge sweet-spot approaches the limit of $2T_{1}$. b, The pure dephasing time $T_{\phi}$ increases as a function of $N$ when the qubit biased away from charge sweet-spot, with a power-law fit of $T_{\phi}\propto N^{0.61}$. c, Repeated Ramsey fringes measured near qubit sweet-spot for 128 iterations, with each record taking 33 s. d, Detuning $\Delta_{dq}$ between drive tone and qubit frequency during the Ramsey measurements, revealing stochastic frequency shifts.
@@
As illustrated in Fig. Noise-resilient solid host for electron qubits above 100 mKc, the local neon profile and the trap structure define the potential energy landscape seen by the electrons. Recent theoretical work emphasized the neon surface’s important role in defining the qubit’s Hamiltonian [12], whose exact form remains unclear. Here, we approximate each eNe qubit’s transition frequency with a generalized hyperbolic model, capturing the measured qubit spectroscopic features:
@@
Bringing eNe qubit’s transition frequency $\omega_{\rm{q}}$ onto resonance with $\omega_{\rm{r}}$ results in the vacuum Rabi splitting in the resonator’s transmission spectrum. An example of this is shown in Fig. Noise-resilient solid host for electron qubits above 100 mKa for Q1, one of the three qubits we characterized, with the electron-photon coupling strength $g/2\pi=6.43$ MHz and on-resonance qubit linewidth $\gamma/2\pi=3.81$ MHz (see Supplementary III-VI for details of characterizations on the three qubits Q1-3). Compared to the previous Nb resonator [6, 7], the higher impedance of the TiN resonator enhances the qubit-resonator coupling strength, with a maximum observed $g/2\pi$ of approximately $16$ MHz (see Q3 in Supplementary VI). We further mapped the qubit spectrum by applying a second drive tone and probing at the resonator frequency at low power, as shown in Fig. Noise-resilient solid host for electron qubits above 100 mKb. The extracted qubit frequency follows a hyperbolic dependence on the DC voltage $V_{\text{res}}$ applied to the resonator with a charge sweet-spot at 5.065 GHz (see Supplementary IV). The sweet-spot frequency, coupling strength, and sensitivity to DC biases ($\partial f_{\text{q}}/\partial V_{\text{res}}$) vary between qubits (see Supplementary III). This variation suggests some randomness in the local trapping potential, which determines the qubit’s minimum transition frequency and the projection of the electron’s dipole moment along the resonator mode and the applied DC field.
@@
Next, we characterize the coherence performance of the qubit biased at its charge sweet-spot. Figure Noise-resilient solid host for electron qubits above 100 mKc-d show the measured $T_{1}$, $T_{2}^{*}$, and $T_{2}^{\text{echo}}$ of 11.6 µs, 8.2 µs, and 22.6 µs, respectively. Unlike the qubit reported in our previous works [7], the relaxation time ($T_{1}$) of eNe qubits observed on the new TiN device is generally not Purcell-limited at their charge sweet-spots, with over 100 MHz detuning from the resonator ($\Delta_{\rm{rq}}/2\pi$). The Purcell rate of Q1 to the resonator mode is $\Gamma_{\rm{r}}\,=\,\kappa g^{2}/\Delta_{\rm{rq}}^{2}\,=\,1/3.9\,\text{ms}^{-1}$, which indicates that non-radiative decay channels dominate the energy relaxation of Q1. Measurements of the Q1’s $T_{1}$ at various bias frequencies (see Noise Spectroscopy section) further confirm that the non-radiative decay dominates the energy relaxation unless the qubit frequency is tuned much closer to the resonator frequency. Additionally, the fact that the $T_{2}^{\rm{echo}}$ approaches $2T_{1}$ indicates that the quasi-static noise is the dominating dephasing factor for Q1 at its charge sweet-spot [37, 38].
@@
Fig. 4. : Noise spectroscopy of eNe qubit Q1. Main plot between 0.01 $\sim$ 1 MHz: Total noise density (colored dots) derived from dynamical decoupling data at different qubit bias points. Main plot between $10^{-3}$ to $10^{-1}$ Hz: Calculated total noise density (blue dots) from long-term Ramsey measurements when biased at charge sweet-spot. Main plot near 5 GHz: Transverse noise of the eNe qubit (green triangles). Inset plot i: Equivalent charge (voltage) noise on the resonator electrode. Inset plot ii: Zoom-in of transverse noise. Gray dashed lines: Power-law fits of frequency-dependent noise.
@@
Fluctuations in the charge environment can cause stochastic frequency shifts of the eNe qubit, which leads to qubit dephasing. Armed with high fidelity control and a known voltage (charge) lever arm, we first characterize the high-frequency noise spectral density of Q1 using the Carr-Purcell-Meiboom-Gill (CPMG) sequence with $N=0,1,2,4,6,8,12$, and 16 refocusing pulses. All bias points had positive $\Delta V_{\rm{res}}$ as shown in Fig. Noise-resilient solid host for electron qubits above 100 mKb, with a maximum sensitivity of $\partial f_{\rm{q}}/\partial V_{\rm{res}}=180.7$ MHz/mV, corresponding to a qubit frequency detune of 15.9 MHz from the charge sweet-spot. The measured average energy relaxation time at all bias points is 11.94 $\pm$ 0.3 µs, indicating that the same non-radiative decay channel dominates. We fit the measured coherence decay to extract the decoherence time $T_{2}^{\text{CPMG}}$ and pure dephasing time $T_{\phi}$ (See Methods). As shown in Fig. Noise-resilient solid host for electron qubits above 100 mKa, in the absence of refocusing pulses (i.e., the Ramsey measurement), the decoherence rate $1/T_{2}^{\text{CPMG}}$ increases with sensitivity to charge noise. At $\partial f_{\rm{q}}/\partial V_{\rm{res}}=180.7$ MHz/mV, $T_{2}^{*}$ decreases to 1.93 µs. Introducing refocusing pulses enhances $T^{\text{CPMG}}_{2}$ at every biasing point, approaching two times of $T_{1}$ as $N$ increased. At the charge sweet-spot, the rapid saturation of $T_{2}^{\text{CPMG}}$ with increasing $N$ indicates a low noise power density in the relatively high-frequency range (100 kHz). When the qubit was biased at a point more sensitive to charge noise, more pulses were required to extend the $T_{2}^{\text{CPMG}}$ towards the energy relaxation limit. In Supplementary V and VI, we show that for the other two eNe qubits, a single refocusing pulse is insufficient to mitigate the majority of noises at the qubits’ sweet-spot, suggesting a higher noise density at the high-frequency range. This variation in noise behavior suggests some complexities of the local noise environment experienced by individual electrons on the neon surface.
@@
To benchmark different electron qubit material platforms, we compare the environmental charge noise measured as voltage fluctuations on adjacent gate electrodes, situated at a distance on the scale of one hundred nanometers from the qubits. Specifically, we examine electrons trapped on solid neon and those confined in semiconductor materials, which are the leading platforms for hosting electron qubits. Table 1 summarizes our results on eNe and compares them with data from the literature on semiconductor qubits. Because of the clean qubit environment provided by the neon interface, the charge noise experienced by eNe is orders of magnitude lower than that in typical semiconductor materials, approaching some of the best performance. Given the smaller susceptibility of electron spin qubits to charge noise through effects including spin-orbit coupling and exchange interactions, as shown in the “$\partial f_{q}/\partial V$” column of Tab. Noise-resilient solid host for electron qubits above 100 mK, we foresee that electron spin qubits on solid neon could approach a pure dephasing time of millisecond-scale [9]. The development of the electron trap and the suppression of redundant electrons on neon will further improve the qubit performance.
@@
After investigating the high-frequency noise, we turn to the low-frequency noise at the charge sweet-spot. Figure Noise-resilient solid host for electron qubits above 100 mKc-d show variations in Q1’s frequency revealed by performing repeated Ramsey measurements [21, 20] for approximately one hour since initially biased near its charge sweet-spot. The qubit undergoes a discrete frequency transition of varying magnitudes, as seen near the 25 minutes of the one-hour tracing. This is revealed by a Fast Fourier transform (FFT) of the Ramsey signals (See Methods and Supplementary IV). It suggests slow mechanisms driving qubit decoherence, akin to low-frequency dynamics in semiconductor and superconducting qubits [21, 20, 40], which can be mitigated with feedback control [41, 42]. Using a periodogram method, we convert this one-hour qubit frequency measurement into a frequency noise spectrum over $10^{-3}$ to $10^{-1}$ Hz, as shown in Fig. Noise-resilient solid host for electron qubits above 100 mK. A power-law fit of the data gives the relation of $S\propto f^{-1.11}$. Despite the qubit being first-order insensitive to small charge (voltage) fluctuation at the sweet-spot, it can still experience slow frequency drifts due to nearby charge fluctuations, as observed in both semiconductor and superconducting qubits via second-order effects [43, 44]. Other detection methods, such as single-electron charge sensing techniques [21], could complete the noise spectrum ranging from 1 Hz to $10^{4}$ Hz.
@@
To complete the discussion, we calculate the transverse noise at the qubit frequencies contributing to energy relaxation with $1/T_{1}=\pi/2\times S(2\pi f_{\rm{q}})$ (ref. [37]). The fitted total transverse noise versus frequency follows $S\propto f^{2.22}$ between 5.065 to 5.498 GHz, with a mean of $3.6\times 10^{5}$$^{2}/\text{Hz}$ Hz, as shown in the inset of Fig. Noise-resilient solid host for electron qubits above 100 mK.
@@
To obtain more information about the noise environment of eNe qubits, we measured the temperature dependence of Q1’s coherence ($T_{1}$, $T_{2}^{*}$, and $T_{2}^{\rm{echo}}$) at its charge sweet-spot from 10 mK to 500 mK. The energy relaxation $T_{1}$ is well described by a model that only accounts for the coupling of a two-level quantum system to a bosonic thermal bath within the experiment’s temperature range, $T_{1}\propto\text{tanh}(\hbar\omega_{\rm{q}}/2k_{\text{B}}T)$ (Fig. Noise-resilient solid host for electron qubits above 100 mKa) (ref. [45, 46]). Similarly, the thermal population follows that of a Maxwell-Boltzmann distribution [47] (See Supplementary IV). The corresponding single electron temperature closely tracks the mixing-chamber (MXC) temperature between 100 to 500 mK and saturates around 40 mK, as shown in Fig. Noise-resilient solid host for electron qubits above 100 mKa. This effective cooling of electron qubits below 100 mK is difficult to achieve with electrons in semiconductors, due to the suppressed electron-environment interaction at low temperatures [48].